Home‎ > ‎Crew 135‎ > ‎

206-CommanderReport

Commander Report
02/06/2014
Ondrej Doule


Today, the Crew 135 started research on structural properties of MDRS. This simulated research
Mars desert storm load may be repeated next week to acquire more data.

MDRS habitat technical sketch is completed. Attached in the mail version. It will be developed and
used for model generation and for the architectural concept of the base expansion.

Greenhab research included also planting of carrots, radish and peas as requested. Future  crews
will have lots of fresh veggies to harvest.

The EVA04 around the habitat was successful. Still the biggest problem are scratched helmets that
do not allow easy reading of data from instruments / precision work. The overall suit set up is
otherwise fine and using red glasses provides really spectacular experience and real feel of Mars.
EVA04 crewmembers we quite tired after return and half of them needed immediate rest.

EVA05 will include Glass experiment and further surveying in the vicinity of the base to the North.

Nutrition experiment is giving interesting unexpected psychological results - some crewmembers eat
lot of uniform food and avoid eating divers and often to avoid repeated nutrition reporting.

Our reporters are doing spectacular job - working non-stop on MDRS and Crew135 PR.

==

Dear Sri,

We are all online, research is going well, Mars is spectacular...

CommanderCheckIn Report
02/06/2014
Ondrej Doule
Crew Physical Health:
All are well. One crewmember scheduled for today EVA04 reported
light cold and fatigue from yesterday EVA03 and therefore was
advised to stay at the base and not participate EVA04.

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:
- Crewmember LP was exchanged for Crewmember FK - fatigue reasons
- Structural experiment performed successfully, though lots dexterity based difficulties was encountered.
- The vicinity of the base was surveyed and measured.

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Hab Maint:

Rep Sched:
   19:00 - 21:00
Tomorrow Plan:
   - Analysis of results and archiving of data (EXP01 - EXP02), preparation of EXP04 - all crew.
- EVA05 - Additional base vicinity survey and HUD experiment starting 16:00 - finished at 18:00 - 4 crewmembers

Inventory:

Support Req:
   XL jumpsuit is requested. Commander Doule (6.3 feet, 193 cm)has hard time fitting in the L size - thanks for understanding
Miscelleneous: